OpenStudy (anonymous):

I need help solving this equation? the blade on a typical table saw rotates at 3000 revolutions per minute. Calculate the linear velocity in feet per second at the edge of a 1 foot diameter blade.

OpenStudy (tkhunny):

\(Circumference = 2\pi r = \pi d\)

OpenStudy (anonymous):

let me know if i'm doing it right okay. i did 3000(3.14)/60 and i got 157.07 which rounds to 157.1 is this correct?

OpenStudy (tkhunny):

Let's see: \(\dfrac{3000\;Rev}{Min}\cdot \dfrac{\pi \cdot 1 ft}{Rev}\cdot \dfrac{1\;min}{60\;sec} = \dfrac{3000\pi}{60}\cdot\dfrac{ft}{sec} = 157.0796\;fps\) I think you have it!! Good work.

OpenStudy (anonymous):

thank you